OpenStudy (anonymous):

Betty bought a package of meat that weighed 2 1/2 pounds and another package that weighed 1 3/4 pounds. how much meat did she buy altogether?

OpenStudy (anonymous):

4 1/4

OpenStudy (anonymous):

1/2 is 2/4 so 2 2/4+ 1 3/4 = 4 1/4

OpenStudy (anonymous):

...

OpenStudy (anonymous):

I just put it in & Plato said it was incorrect.

OpenStudy (anonymous):

2&2/4+1&3/4=3&5/4=4&1/4